# Opponent Shaping for Antibody Development

### v4

@@ 3 Discussion
This study introduces a novel framework for designing antibodies using the principle of opponent shaping [9] that can effectively combat evolving pathogens. The core of our framework is a two-player game that models the interaction between antibodies and viruses. Applying opponent shaping to the game allows us to optimise for antibody shapers that anticipate and influence viral evolution.
@@
The proposed framework is adaptable. While we demonstrated its efficacy using the Absolut! [11] binding simulator accelerated with JAX [13] for payoff estimation, and our model of the simulated viral escape with an evolutionary process, these components are interchangeable. As more accurate and efficient binding simulators or viral evolution models become available, they can be integrated into our framework, potentially enhancing its biological applicability.
@@
Our results demonstrate a clear advantage of antibody shapers over myopic antibodies when considering long horizons of viral escape. We observed that myopic antibodies, while initially effective, struggle to maintain their efficacy as the virus evolves. In contrast, shaper antibodies, particularly those optimized for longer horizons, show sustained performance over extended periods of viral evolution, see Figure 2.
A notable computational trade-off emerges in the performance of antibody shapers. While longer optimisation horizons generally lead to better long-term performance, they come at an increased computational cost. Figure 2d offers guidance for selecting optimal optimisation horizons based on available computational resources and runtime of a binding simulation which is particularly valuable for future applications of our framework.
@@
Furthermore, our analysis reveals that long-horizon shapers shape the virus into more exploitable variants as opposed to just becoming robust to any virus, see Figure 3. In fact, the short-horizon shapers bind stronger to the viruses included by the long-horizon shapers than the long-horizon shapers themselves. This suggests exploring a mixture of antibodies, some optimised for virus shaping and others for binding. This could leverage the benefits of both shaping the viral evolutionary landscape and maintaining strong binding efficacy. Interestingly we observe a similar “strategy” deployed by the human body: a longitudinal study on rare patients who naturally developed broadly neutralizing antibodies (bnAbs) to HIV showed that early neutralizing antibodies induce HIV escape to those antibodies [23]. In the future, our approach could enable us to better understand and develop therapies for similar settings by simulating a game with more than one antibody playing against the virus.
@@
To provide insights into the mechanisms underlying the effectiveness of antibody shapers, we conduct an explainability analysis focused on the amino acid distribution in shapers (Figure 4) and on binding poses (Figure 5). While these results offer insights into the strategies employed by our antibody shapers, it’s important to note that they are specific to the Absolut! binding model [11, 14], and the transferability of our findings to more realistic binding simulators and viral escape models remains an open question. Antibodies designed using more accurate binding simulators might be able to employ other strategies in their interactions with viruses, and the strategies we explore in this model may not apply in a more realistic simulation. Nevertheless, the fact that we can explain the behaviour of our antibody shapers is encouraging. It suggests that, even as we move towards more complex and realistic models, we may be able to maintain a degree of interpretability in our approach.
@@
A significant limitation of our current approach is the static structure of the viral antigen. In reality, mutations to the virus induce changes in the structure of the viral antigen. However, in our experiments, we consistently use the original structure of the 2R29 dengue viral antigen [22] and a fixed set of Absolut!-generated docking poses based on that structure. As we introduce mutations to both the virus and the antibody, the identity of the lowest-energy pose changes, but the underlying set of possible poses remains constant. This simplification potentially overlooks important structural dynamics that influence antibody-antigen interactions. Ideally, we could model the structural changes that occur with each introduced mutation to capture a more realistic representation of the evolving virus-antibody interaction landscape. AlphaFold3 has demonstrated the capability to predict structures of antibody-antigen complexes from their sequences with high accuracy [24]. Incorporating methods like AlphaFold3 into our framework would allow for the dynamic updating of antigen structures in response to our simulated mutations. This would enhance the biological relevance of our method.
@@ 4 Methods
Our method applies the principles of opponent shaping [9] to a two-player game between the antibody shaper player or agent and a naive virus agent (Figure 1a). In this section, we describe our method in detail.
Firstly, we introduce the virus-antibody game in Section 4.1 where we define the action spaces and the payoffs of both players. Secondly, we define the simulated viral evolution process in Section 4.2 where the virus evolves to escape in response to the current antibody. Finally, in Section 4.3 we define the antibody optimisation process where we optimise the antibody shapers in a way that accounts for future viral mutations and learns to influence viral evolution away from escape.
Importantly, implementing the virus-antibody game requires a way to determine the strength of binding of any antibody-antigen pair, or as we refer to it later in this section a binding function. In section 4.4 we describe our binding function choice, based on the Absolut! framework [11].
@@ 4.1 Virus-Antibody Game
With $R_{v}(v,a)=-R_{a}(v,a)$. In this framework, the antibody gets a higher payoff not only if it binds the virus with higher affinity, but also if the virus fails to bind its receptor or if the antibody has a lower risk of autoimmunity.
We can now define the simulated viral escape via evolution. In the following section, we leverage the viral payoff $R_{v}(v,a)$ as a fitness function to model how viruses might evolve to evade antibody binding. This simulates viral trajectories that viruses could take in response to therapeutic pressure.
@@ 4.2 Simulated Viral Escape via Evolution
At generation $i$ we have a population of $P$ viruses $v^{i}_{1},v^{i}_{2}\dots v^{i}_{P}$. In our experiments $P=15$. When $i=0$, the population is the starting virus $\hat{v}$ duplicated $P$ times. For every virus in the population, we evaluate its fitness given by $R_{v}(v^{i}_{k},a).$ We then sample a new virus $\hat{v}^{i}$ based on the fitness values, in particular:
$\mathbb{P}(\hat{v}^{i}=v^{i}_{k})\propto\exp(\beta R_{\text{v}}(v^{i}_{k},a))$
With duplicates in the population being considered distinct, the likelihood of a particular variant increases with the number of duplicates. Furthermore, $\beta$ is a constant which reflects how random the selection process is, with $\beta\rightarrow\infty$ reflecting deterministic max-fitness selection.
To generate the next population, $\hat{v}^{i}$ is then duplicated and randomly mutated $P$ times such that on average there is one amino acid mutation in the viral sequence, so:
$v^{i+1}_{k}=\hat{v}^{i}\oplus\textit{Mutation}$
After horizon $H$ generations, a full escape trajectory $\bm{\hat{v}}=\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}$ is generated and the simulated viral escape process ends.
@@ 4.3 Antibody Optimisation
We define the antibody fitness $F^{H}_{\hat{v}}(a)$ such that it represents the true objective of the antibody which accounts for the viral escape. Given a horizon $H$ and starting virus $\hat{v}$, antibody fitness is:
@@
This is a precise definition of the antibody fitness, Equation 1 in Section 2.1.3 uses a simplified notation. Note that if $H=0$ this fitness defaults to a naive antibody payoff that ignores viral escape, i.e. $F^{0}_{\hat{v}}(a)=R_{a}(\hat{v},a)$. We refer to this as the myopic objective.
To optimise antibody shapers, we employ Monte Carlo simulations to estimate the antibody fitness combined with an evolutionary algorithm. We refer to this process as the antibody optimisation loop, see Figure 1a. In meta-learning terms, this is the outer loop or the meta-loop, contrasting to the inner loop which is the simulated viral escape via evolution.
@@
At antibody optimisation step $i>0$, we have a population of $P_{a}$ antibodies $a^{i}_{1},a^{i}_{2}\dots a^{i}_{P_{a}}$. In our experiments $P_{a}=40$. We then sample their fitness values $F^{H}_{\hat{v}}(a^{i}_{j})$ with a fixed number $\eta$ of Monte Carlo roll-outs, i.e. we sample $\eta$ independent viral escape trajectories each with horizon $H$ viral escape steps. We found $\eta=5$ to be sufficient. We then select the best-performing antibody, $\hat{a}^{i}$:
$\hat{a}^{i}=\arg\max_{k}\mathbb{E}\left[F_{\hat{v}}^{H}(a^{i}_{k})\right]$
When $i=0$, $\hat{a}^{0}=\hat{a}$ is the starting random antibody. We generate the next population of antibodies $a^{i+1}_{1},a^{i+1}_{2}\dots a^{i+1}_{P_{a}}$ by taking both the antibody $\hat{a}^{i}$ and $P_{a}-1$ copies of it, each copy with a single random mutation to the amino acid sequence of $\hat{a}^{i}$.
Once the final optimised antibody $\hat{a}_{N}$ is generated, a full optimisation trajectory is complete $\bm{\hat{a}}=\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}$ and the antibody optimisation process finishes.
@@ 4.4 Binding Function
The energy of a complex of a virus $v\in\mathbb{A}^{N_{v}}$ and an antibody $a\in\mathbb{A}^{N_{a}}$, in a given pose $(\phi^{v\times a},\phi^{a\times a})\in\Phi$ is defined by sum of the energies of each adjacent residues. The energy between a residue pair is determined by which two amino acids it contains, given by a symmetric interaction matrix $M:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{R}$ which is determined experimentally by [14].
@@
To give a more concrete example, for this paper, we use the dengue virus antigen [22]. Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation 1000100010001000555In practice, the difference is closer to $10,000$ times faster for a negligible performance drop for this particular antigen sequence.
However, this leads to more errors as soon as we change the viral antigen sequence. Looking at the particular poses which lead to binding does reveal another way to cut down on the total number of poses: all of the poses contain at least 18 pairs of residues. As the interaction matrix $M$ is strictly negative, having more pairs of residues always makes the binding energy of a pose lower, meaning it is more likely to be where binding occurs. Out of the original 1.5 million poses, only approximately 37 thousand (1 in 40) contain 18 or more pairs of residues. When using only these residues, we see no differences across any of the evolutionary simulations. It is possible that a pose with 17 or less pairs is the dominant one for some antigen $v$ with antibody $a$, but if so such pairs appear to be extremely rare.
@@
When running experiments, we always train with the low-resolution binding simulator, then perform “verification” with the high-resolution one and these are the results we report throughout the paper. The reason is twofold. Firstly this enables us to run many more evolution experiments. Secondly, this mimics the real-life process of transferring out of simulation to the real world. By showing we transfer from the low-resolution binding simulation to the slower, high-resolution binding simulation, we demonstrate that our results are not extremely specific to the exact simulation we use and that any result will not disappear as soon as a more accurate simulation is used.
We emphasise that Absolut! does not represent an accurate model of antibody binding. It is instead a toy simulation to demonstrate our methodology. For example, we do not expect our framework when used with this simulation model to yield superior, effective antibodies.

### v6

@@ 3 Discussion
This study introduces a novel framework for designing antibodies using the principles of opponent shaping [9] that can effectively combat evolving pathogens. The core of our framework is a two-player game that models the interaction between antibodies and viruses. Applying opponent shaping to the game allows us to optimise for antibody shapers that anticipate and influence viral evolution.
@@
Our proposed framework is adaptable. While we demonstrated its efficacy using the Absolut! [11] binding simulator accelerated with JAX [13] for payoff estimation, and our model of the simulated viral escape with an evolutionary process, these components are interchangeable. As more accurate and efficient binding simulators or viral evolution models become available, they can be integrated into our framework, potentially enhancing its biological applicability.
@@
Our results demonstrate a clear advantage of antibody shapers over myopic antibodies when considering long horizons of viral escape. We observed that myopic antibodies, while initially effective, struggle to maintain their efficacy as the virus evolves. In contrast, shaper antibodies, particularly those optimized for longer horizons, show sustained performance over extended periods of viral evolution, see Figure 2. A notable computational trade-off emerges in the performance of antibody shapers. While longer optimisation horizons generally lead to better long-term performance, they come at an increased computational cost. Figure 2d offers guidance for selecting optimal optimisation horizons based on available computational resources and runtime of a binding simulation which is particularly valuable for future applications of our framework.
@@
Furthermore, our analysis reveals that long-horizon shapers shape the virus into more exploitable variants as opposed to just becoming robust to any virus, see Figure 3. In fact, the short-horizon shapers bind stronger to the viruses induced by the long-horizon shapers than the long-horizon shapers themselves. This suggests exploring a mixture of antibodies, some optimised for virus shaping and others for binding, leveraging the benefits of both shaping the viral evolutionary landscape and maintaining strong binding efficacy. Interestingly we observe a similar “strategy” deployed by the human body: a longitudinal study on rare patients who naturally developed broadly neutralizing antibodies (bnAbs) to HIV showed that early neutralizing antibodies induce HIV escape to those antibodies, resulting in enhanced binding to other antibodies [23]. In the future, our approach could enable us to better understand and develop therapies for similar settings by simulating a game with more than one antibody playing against the virus, drawing inspiration from broader research on coevolving systems and their complex interactions [24, 25].
@@
To better understand the mechanisms underlying the effectiveness of antibody shapers, we conduct an explainability analysis focused on the amino acid distribution in shapers (Figure 4) and on binding poses (Figure 5). While these results offer insights into the strategies employed by our antibody shapers, it’s important to note that they are specific to the Absolut! binding model [11, 14], and the transferability of our findings to more realistic binding simulators and viral escape models remains an open question. Antibodies designed using more accurate binding simulators might be able to employ other strategies in their interactions with viruses, and the strategies we explore in this model may not apply in a more realistic simulation. Nevertheless, the fact that we can explain the behaviour of our antibody shapers is encouraging. It suggests that, even as we move towards more complex and realistic models, we may be able to maintain a degree of interpretability in our approach.
@@
A significant limitation of our current approach is the static structure of the viral antigen. In reality, mutations to the virus induce changes in the structure of the viral antigen. However, in our experiments, we consistently use the original structure of the 2R29 dengue viral antigen [22] and a fixed set of Absolut!-generated docking poses based on that structure. As we introduce mutations to both the virus and the antibody, the identity of the lowest-energy pose changes, but the underlying set of possible poses remains constant. This simplification potentially overlooks important structural dynamics that influence antibody-antigen interactions. Ideally, we could model the structural changes that occur with each introduced mutation to capture a more realistic representation of the evolving virus-antibody interaction landscape. AlphaFold3 has demonstrated the capability to predict structures of antibody-antigen complexes from their sequences with high accuracy [26]. Incorporating methods like AlphaFold3 into our framework would allow for the dynamic updating of antigen structures in response to our simulated mutations. This could enhance the biological relevance of our method.
@@ 4 Methods
Our method applies the principles of opponent shaping [9] to a two-player game between the antibody shaper player or agent and a naive virus agent (Figure 1a). In this section, we describe our method in detail. Firstly, we introduce the virus-antibody game in Section 4.1 where we define the action spaces and the payoffs of both players. Secondly, we define the simulated viral evolution process in Section 4.2 where the virus evolves to escape in response to the current antibody. Finally, in Section 4.3 we define the antibody optimisation process where we optimise the antibody shapers in a way that accounts for future viral mutations and learns to influence viral evolution away from escape. Importantly, implementing the virus-antibody game requires a way to determine the strength of binding of any antibody-antigen pair, or as we refer to it later in this section a binding function. In Section 4.4 we describe our binding function choice, based on the Absolut! framework [11].
@@ 4.1 Virus-Antibody Game
With $R_{v}(v,a)=-R_{a}(v,a)$. In this framework, the antibody gets a higher payoff not only if it binds the virus with higher affinity, but also if the virus fails to bind its receptor or if the antibody has a lower risk of autoimmunity. We can now define the simulated viral escape via evolution. In the following section, we leverage the viral payoff $R_{v}(v,a)$ as a fitness function to model how viruses might evolve to evade antibody binding. This simulates viral trajectories that viruses could take in response to therapeutic pressure.
@@ 4.2 Simulated Viral Escape via Evolution
At generation $i$, we have a virus $\hat{v}^{i}$. We then generate a population of viruses $v^{i}_{1},v^{i}_{2}\dots v^{i}_{P}$ by duplicating $\hat{v}^{i}$ $P$ times, then randomly applying mutation such that on average there is one amino acid mutation per viral sequence:
$v^{i}_{k}=\hat{v}^{i}\oplus\textit{Mutation}$
In our experiments $P=15$. For every virus in the population, we evaluate its fitness given by $R_{v}(v^{i}_{k},a).$ We then sample a new virus $\hat{v}^{i+1}$ based on the fitness values, in particular:
$\mathbb{P}(\hat{v}^{i+1}=v^{i}_{k})\propto\exp(\beta R_{\text{v}}(v^{i}_{k},a))$
With duplicates in the population being considered distinct, the likelihood of a particular variant increases with the number of duplicates. Furthermore, $\beta$ is a constant which reflects how random the selection process is, with $\beta\rightarrow\infty$ reflecting deterministic max-fitness selection. After $H$ generations, $H$ being our chosen horizon, a full escape trajectory $\bm{\hat{v}}=\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}$ has been generated and the simulated viral escape process ends.
@@ 4.3 Antibody Optimisation
We define the antibody fitness $F^{H}_{\hat{v}}(a)$ such that it represents the true objective of the antibody, which accounts for the viral escape. Given a horizon $H$ and starting virus $\hat{v}$, the antibody fitness is:
@@
As opposed to Equation 2 in Section 2.1.3, which uses a simplified notation, this is a more formal definition of the antibody fitness. Note that if $H=0$ this fitness defaults to a naive antibody payoff that ignores viral escape, i.e. $F^{0}_{\hat{v}}(a)=R_{a}(\hat{v},a)$. We refer to this as the myopic objective.
To optimise antibody shapers, we employ Monte Carlo simulations to estimate the antibody fitness, combined with an evolutionary optimisation algorithm. We refer to this process as the antibody optimisation loop, see Figure 1a. In meta-learning terms, this is the outer loop or the meta-loop, contrasting to the inner loop which is the simulated viral escape via evolution.
@@
At the start of antibody optimisation step $i$, we have an antibody $\hat{a}^{i}$. We first generate a population of $P_{a}$ antibodies $a^{i}_{1},a^{i}_{2}\dots a^{i}_{P_{a}}$ by taking both the antibody $\hat{a}^{i}$ and $P_{a}-1$ copies of it, with each copy having a single random mutation to the amino acid sequence of $\hat{a}^{i}$. For our experiments, $P_{a}=40$.
We then sample their fitness values $F^{H}_{\hat{v}}(a^{i}_{j})$ with a fixed number $\eta$ of Monte Carlo roll-outs, i.e. we sample $\eta$ independent viral escape trajectories each with horizon $H$ viral escape steps. We found $\eta=5$ to be sufficient. Finally, we select $\hat{a}^{i+1}$ to be the best-performing antibody:
$\hat{a}^{i+1}=\arg\max_{k}\mathbb{E}\left[F_{\hat{v}}^{H}(a^{i}_{k})\right]$
Once the final optimised antibody $\hat{a}^{N}$ is generated, a full optimisation trajectory is complete, $\bm{\hat{a}}=\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}$, and the antibody optimisation process finishes.
@@ 4.4 Binding Function
The energy of a complex of a virus $v\in\mathbb{A}^{N_{v}}$ and an antibody $a\in\mathbb{A}^{N_{a}}$, in a given pose $(\phi^{v\times a},\phi^{a\times a})\in\Phi$ is defined by sum of the energies of each adjacent residues. The energy between a residue pair is determined by which two amino acids it contains, given by a symmetric interaction matrix $M:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{R}$, which is determined experimentally [14].
@@
To give a more concrete example, for this paper, we use the dengue virus antigen [22]. Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation $1000$ times faster [^5] In practice, the difference is closer to $10,000$, likely due to the GPU having to move less data. for a negligible accuracy drop for this particular antigen sequence.
However, this leads to more errors as soon as we change the viral antigen sequence. Looking at the particular poses which lead to binding does reveal another way to cut down on the total number of poses: all of the poses contain at least 18 pairs of residues. As the interaction matrix $M$ is strictly negative, having more pairs of residues always makes the binding energy of a pose lower, meaning it is more likely to be where binding occurs. Out of the original 1.5 million poses, only approximately 37 thousand (1 in 40) contain 18 or more pairs of residues. When using only these residues, we see no differences across any of the evolutionary simulations. It is possible that a pose with 17 or less pairs is the dominant one for some antigen $v$ with antibody $a$, but if so then such pairs appear to be extremely rare.
@@
When running experiments, we always train with the low-resolution binding simulator, then perform “verification” with the high-resolution one and these are the results we report throughout the paper. The reason is twofold. Firstly this enables us to run many more evolution experiments. Secondly, this mimics the real-life process of transferring out of simulation to the real world. By showing we transfer from the low-resolution binding simulation to the slower, high-resolution binding simulation, we demonstrate that our results are not extremely specific to the exact simulation we use and that any result will not disappear as soon as a more accurate simulation is used. We emphasise that Absolut! does not represent an accurate model of antibody binding. It is instead a toy simulation to demonstrate our methodology. For example, we do not expect our framework when used with this simulation model to yield highly effective, superior antibodies for real-world applications.